Find the equation of the line in slope-intercept form that is perpendicular to y=-1/4x+3 and passes through (1.5,5).
Answer by Cromlix(4381)

Lines that are perpendicular to
one another have slopes that
multiply together to give -1
The line that is perpendicular to
y = -1/4x + 3 has a slope = 4.
-1/4 * 4 = -1
Using the equation:
y - b = m(x - a)
a = 1.5, b = 5 and m = 4
y - 5 = 4(x - 1.5)
y - 5 = 4x - 6
y = 4x - 6 + 5
y = 4x - 1
Hope this helps.
:-)